# MultiTok: Variable-Length Tokenization for Efficient LLMs Adapted from LZW Compression

### v1

@@ Abstract
Large language models have drastically changed the prospects of AI by introducing technologies for more complex natural language processing. However, current methodologies to train such LLMs require extensive resources including but not limited to large amounts of data, expensive machinery, and lengthy training. To solve this problem, this paper proposes a new tokenization method inspired by universal Lempel-Ziv-Welch data compression that compresses repetitive phrases into multi-word tokens. With MultiTok as a new tokenizing tool, we show that language models are able to be trained notably more efficiently while offering a similar accuracy on more succinct and compressed training data. In fact, our results demonstrate that MultiTok achieves a comparable performance to the BERT standard as a tokenizer while also providing close to 2.5x faster training with more than 30% less training data.
@@ I Introduction
Additionally, methods such as transferring knowledge across language models[6] and creating sparse representation for data [7] have been utilized to reduce the training cost. These methods utilize post-processing techniques to reduce model parameters. However, many of these solutions are not dynamic and still require some baseline tokenization or rely on specific linear data distributions. There also exists methods to reduce inference cost via removing unnecessary neural parameters through quantization[8]. Still, these methodologies are not as effective since they are post processing methods, require a large amount of resources, and are computationally costly to run. A promising direction includes using compression algorithms on training data to expedite the training [9]. One of the most popular lossless compression algorithms is the Lempel-Ziv-Welch (LZW) algorithm [10]. The LZW algorithm is a universal compression algorithm that works by compressing repeated patterns of bits within a dataset. As such, it achieves a high compression ratio and performs better with larger amounts of data [11].
In this paper, we introduce MultiTok, a new tokenization method inspired by the LZW compression algorithm for the efficient LLM training tasks. We focus on LLMs where the inputted data is first tokenized utilizing an encoder, extracted into embedding vectors representing the semantic and contextual information, and finally fed into a transformer model for text classification. MultiTok provides a novel variable-length tokenizer in the sense that each token can represent a variable number of words. The advantages of MultiTok are it (i) dynamically compresses the necessary training data by close to $33\%$ (ii) allows the LLM model to be trained close to 2.5x times faster and (iii) maintains performance comparable to the baseline, e.g., Bidirectional Encoder Representations from Transformers (BERT) [12] standard. Our tokenizing schemes can mark the beginning of the use of information-theoretic approaches to provide efficient, secure, and robust LLM systems. The code and data can be accessed at: https://github.com/noelkelias/multitok.
@@ II Problem Setting
$x=\{x_{1},\dots,x_{n(x)}\}\rightarrow T(x)=\{y_{1},\dots,y_{m(x)}\}.$ (2)
@@
We consider several performance goals: The utility goal is to train a competitive downstream classification with a high generalization accuracy using the proposed tokenization scheme. The utility is often quantified via the generalization accuracy over some validation data. The efficiency goal is two-fold: compression (having fewer tokens per sample) and convergence (reducing the training time).
@@ III MultiTok: A Variable-Length Tokenization
1:$\mathcal{X},D,w$ 2:$\mathcal{T}=\{T(x)\}_{x\in\mathcal{X}}$ 3:$\mathcal{T}=\{\}$ 4:for $x\in\mathcal{X}$ do 5:   $i=1,T(x)=[{}],idx=|\mathcal{D}|+1,$ 6:   while $i\leq|x|$ do 7:      for $j\in[i,\text{min}(i+w,|x|)]$ do 8:         if $x[i:j]\notin D$ then 9:           $D(x[i:j])=idx,\;idx=idx+1$ 10:           $T(x){+\!=}D(x[i:j-1])\;i:=j$ 11:           break           12:         if $j=|x|$ then $T(x){+\!=}D(x[i:j])$                 13:      $\mathcal{T}=\mathcal{T}\cup T(x)$
@@
We also propose a post-processing technique, where we conduct frequency analysis on the MultiTok dictionary. Particularly, multi-word tokens that only appear a few number of times within the encoded training sample are pruned. As a result, during the encoding, those larger tokens are replaced by their original single-word tokens. The motivation for this step is to provide a smaller dictionary of more frequently utilized tokens for a training task rather than including multi-word tokens that might have been rarely used during tokenization.
Finally, we present the end-to-end tokenization pipeline utilizing Algorithm 1 through an example in Fig. 1. We begin by tokenizing the phrase “Alice goes to the Wonderland and Bob goes to the Wonderland Zoo”. We first find the largest existing phrase in the inputted dictionary (left) and thus tokenize “Alice” as 2. Then, we add “Alice goes” to the MultiTok dictionary (right) as a new token with index 10. We continue this process until the second mention of the phrase “goes to”. This phrase is already in the MultiTok dictionary and is thus encoded as 10 instead of tokens 1 and 5. Now, we add “goes to the” to the MultiTok dictionary with index 16. Afterwards, the tokenized samples undergo frequency analysis (purple). Particularly, we prune tokens like token 12 which were only used twice within the entire training data set $\mathcal{X}$. Thus, its instances in the dataset are replaced with the two tokens that it compresses, i.e., 7 and 4. The resulting tokenized sequence is inputted into the LLM for training.
@@ IV Simulations Results
In this section, we demonstrate the empirical performance of our tokenizer, MultiTok, with respect to the utility and efficiency goals described by Definitions 1 and 2. Further, we show various goals such as utility, accuracy, and training time can compete, and we demonstrate some parameters that can offer a trade-off among them.
@@ IV-A Experimental Setup
The general pipeline includes tokenizing the training samples with the MultiTok and BERT tokenization schemes. These tokens were then assigned corresponding random vector embedding and inputted to train a selected model for the textual binary classification task.
Training. The datasets utilized in this section are the IMDB [13], the sst2 [14], and the AG-News [15] datasets. The first two are collections of movie reviews, each annotated with a binary sentiment. The third is a set of news pieces from different categories. Our experiments were performed on Google Colab instances (Xeon Processors @2.3Ghz with 12GB of RAM). The selected model utilizes a random vector embedding of dimension $100$ per token. This is attached to a deep neural network with an LSTM layer and two hidden linear layers with dimensions $100$ and $50$, respectively. Finally, these nodes are flattened to an output layer with the binary classification output. All models were trained for $30$ epochs utilizing batch sizes of $1000$, the Adam’s optimizer, binary cross-entropy loss, and a learning rate of $0.01$. The complete model configuration and hyperparameters can be found in our code repository https://github.com/noelkelias/multitok.
Evaluation. All our results are the averages over 20 trials with the same parameters. Specifically, we record the model’s average training loss, training accuracy, and testing accuracy per epoch. In addition, the compression ratio as well as the average area under the curve (AUC) given by the receiver operating characteristic curve (ROC) for each test were recorded and graphed. Our results for MultiTok were compared with BERT tokenization [12] and single-word tokenization where each unique word is considered as a token. For all these schemes, we use random vectors for embedding of each token.
Our first round of experiments are focused on studying the performance MultiTok algorithm for various sizes of the training window and the testing window. In the second experiment, we apply MultiTok encoding on the BERT tokens to have a cascade of two tokenization schemes, where MultiTok takes BERT tokens as units of data for compression. Finally, we apply the frequency-based post processing on the MultiTok Tokens for further improvements. We note that the title of rows in the table can be read as the type of tokenization, the percentage of training data to which the tokenizer is applied, and the training and testing window sizes.
@@
Figure 3: MultiTok demonstrates an average training time ($C(\epsilon)|\epsilon=0.01$) that is almost 2-3 times faster than the BERT.
@@
BERT 1.1 0.739 0.81
No Compression 1 0.732 0.81
MultiTok, 100%, Max-Max 0.57 0.644 0.68
MultiTok, 100%, Max-2 0.57 0.637 0.71
@@
MultiTok, 100%, 2-2 0.63 0.705 0.76
MultiTok, 50%, Max-Max 0.8 0.659 0.71
MultiTok, 50%, Max-1 0.8 0.739 0.77
MultiTok, 50%, 2-2 0.83 0.704 0.75
MultiTok, 50%, 2-1 0.83 0.747 0.81
MultiTok, 25%, 2-1 0.93 0.713 0.80
BERT-MultiTok, 50%, 2-1 0.75 0.727 0.78
BERT-MultiTok, 25%, 2-1 0.91 0.745 0.82
MultiTok, 100%, 2-2, $\geq$2 0.66 0.724 0.79
MultiTok, 100%, Max-Max, $\geq$2 0.65 0.707 0.76
@@ IV-B MultiTok Tokenization
Table I shows the results for tokenization schemes with different parameters. According to this table, the best performance is obtained by the presented MultiTok tokenization when the size of training and testing windows are $2$ and $1$, respectively, and the tokenization dictionary is applied to $50\%$ of samples in the training data. We remind that the testing window size $1$ means that during the inference from the trained downstream model, we used single-word tokens, and multi-word tokens were only used during training. In this scenario, MultiTok tokenization exhibits a slightly better performance compared to the standard BERT’s performance ($0.747$ compared to $0.739$) on the IMDB dataset, while compressing the training data by $17\%$.
@@
IV-C MultiTok Tokenization on BERT Tokens
In this experiment, we used MultiTok tokenization to further compress the repetitive sequences within BERT tokens, and consequently, reduce the training data size while achieving similar performance to that of normal uncompressed BERT tokens. We observed that we could compressed the training data that are tokenized via BERT by $9\%$ using MultiTok, while preserving the accuracy of the downstream model. However, when we further reduced training data size, a higher compression is obtained at the cost of slight degradation in the performance. This suggests a notion of using a ‘golden ratio’ of MultiTok compression to maximize efficiency and performance. This ‘golden ratio’ may vary based on the repetition and succinctness found within different datasets.
@@ IV-D MultiTok Tokenization with Post-Processing
Our results in this experiment indicates that the post processing of MultiTok Tokenization with frequency analysis component dramatically increases the performance of the resulting tokens. We remind that without this post processing, there is a significant decrease in the model’s performance when the full capacity of MultiTok is utilized (when MultiTok is applied to the entire dataset with maximum window sizes), see Table I. However, by pruning less-utilized tokens, we were able to circumvent any impact to accuracy, as shown by the last set of experimental results. Specifically for experiment “MultiTok, 100%, 2-2, $\geq$2” (MultiTok is applied to the entire dataset with a training window size of 2), we see that by simply removing all multi-word tokens that only appear once in the training data, we get an accuracy of around $72\%$. This is comparable to the BERT standard and reduces the training data size by $34\%$
Next, we provide an empirical training convergence analysis, which shows that MultiTok not only performs as good as the BERT standard in terms of the accuracy and AUC, but also results in a faster training process for the downstream models. According to Fig. 2, using MultiTok tokens results in a comparable accuracy almost $2.5x$ faster than when using BERT tokens. Specifically, at epoch 12 in Fig. 2, the loss of the MultiTok experiments have almost achieved their stable minimum value, being at least 0.3 lower than the training loss for the baselines at the same epoch. In addition, as shown in Fig. 3, the average training time (as defined in Definition  2) for MultiTok is around 12-13 epochs, which is significantly lower than the other counterparts. Significantly, MultiTok has the lowest compression ratio out of the selected experiments, where the MultiTok tokens not only compressed the inputted training data by a factor of around $3$, but also helped the model learn the training data almost $2.5x$ faster. This confirms that with MultiTok tokenization, the model is able to learn the training data faster and more efficiently as opposed to the baseline tokenization schemes.

### v2

@@ Abstract
Large language models have drastically changed the prospects of AI by introducing technologies for more complex natural language processing. However, current methodologies to train such LLMs require extensive resources including but not limited to large amounts of data, expensive machinery, and lengthy training. To solve this problem, this paper proposes a new tokenization method inspired by universal Lempel-Ziv-Welch data compression that compresses repetitive phrases into multi-word tokens. With MultiTok as a new tokenizing tool, we show that language models are able to be trained notably more efficiently while offering a similar accuracy on more succinct and compressed training data. In fact, our results demonstrate that MultiTok achieves a comparable performance to the BERT and GPT-2 standards as both a stand-alone tokenizer and an add-on to existing tokenizers while also providing close to 2.5x faster training with more than 30% less training data.
@@ I Introduction
Additionally, methods such as transferring knowledge across language models[6] and creating sparse representation for data [7] have been utilized to reduce the training cost. These methods utilize post-processing techniques to reduce model parameters. However, many of these solutions are not dynamic and still require some baseline tokenization or rely on specific linear data distributions. There also exists methods to reduce inference cost via removing unnecessary neural parameters through quantization[8]. Still, these methodologies are post processing methods, require a large amount of resources for training, and are computationally costly to run.
Tokenization is an essential component of many current large language pipelines [9, 10, 11], which directly affects the volume of training data and training complexity. Some examples of tokenization algorithms in the literature are variants of the WordPiece [12] and Byte Pair Encoding [13, 14]. These algorithms are building blocks for generating the state-of-the-art tokens utilized in the Bidirectional Encoder Representations from Transformers (BERT) [15] and Generative Pre-trained Transformer 2 (GPT-2) [16] models.
A promising direction for addressing the challenge of large training data size is using compression algorithms on training data [17, 14]. Specifically, compression algorithms like Huffman coding[18] and other lossless compression algorithms[17] have been evaluated as tokenizers[19, 20, 21]. In addition, implementing linear algebraic error-correcting codes have also been in-play [22, 23]. However, these schemes fall short in their performance or their computational cost with respect to traditional tokenizers.
One of the most well-known lossless compression algorithms is the Lempel-Ziv-Welch (LZW) algorithm [24]. The LZW algorithm is a universal compression algorithm that works by compressing repeated patterns of bits within a dataset. As such, it achieves a high compression ratio and performs better with larger amounts of data [25]. In this paper, we introduce MultiTok, a new tokenization method inspired by the LZW compression algorithm for the efficient LLM training tasks. We focus on LLMs where the inputted data is first tokenized utilizing an encoder, extracted into embedding vectors representing the semantic and contextual information, and finally fed into a transformer model for text classification. MultiTok provides a novel variable-length tokenizer in the sense that each token can represent a variable number of words. The advantages of MultiTok are it (i) dynamically compresses the necessary training data by $33\%$ (ii) allows the LLM to be trained close to 2.5x times faster and (iii) maintains performance comparable to the baseline BERT and GPT-2 token standards. Our tokenizing schemes can mark the beginning of the use of information-theoretic approaches to provide efficient, secure, and robust LLM systems. The code and data can be accessed at: https://github.com/noelkelias/multitok.
@@ II Problem Setting
$x=[x_{1},\dots,x_{n(x)}]\rightarrow T(x)=[y_{1},\dots,y_{m(x)}].$ (2)
@@
We consider several performance goals: The utility goal is to train a competitive downstream classifier with a high generalization accuracy using the proposed tokenization scheme. The utility is often quantified via the generalization accuracy over some validation data. The efficiency goal is two-fold: compression (having fewer tokens per sample) and convergence (reducing the training time).
@@ III MultiTok: A Variable-Length Tokenization
1:$\mathcal{X},D,w$ 2:$\mathcal{T}=\{T(x)\}_{x\in\mathcal{X}}$ 3:$\mathcal{T}=\{\}$ 4:$idx=|\mathcal{D}|+1$ 5:for $x\in\mathcal{X}$ do 6:   $i=1,T(x)=[{}]$ 7:   while $i\leq|x|$ do 8:      for $j\in[i,\text{min}(i+w,|x|)]$ do 9:         if $x[i:j]\notin D$ then 10:           $D(x[i:j])=idx,\;idx=idx+1$ 11:           $T(x){+\!=}D(x[i:j-1]),\;i:=j$ 12:           break           13:         if $j=|x|$ then $T(x){+\!=}D(x[i:j])$                 14:      $\mathcal{T}=\mathcal{T}\cup T(x)$
@@
We also propose a post-processing technique, where we conduct frequency analysis on the MultiTok dictionary. Particularly, multi-word tokens that only appear a few number of times within the encoded training sample are pruned. As a result, during the encoding, those larger tokens are replaced by their smaller constituent tokens that are more frequently used. The motivation for this step is to provide a smaller dictionary of more frequently utilized tokens for a training task rather than including multi-word tokens that might have been rarely used during tokenization. According to our experiment (see Table I), by pruning the tokens that are utilized less than twice both the accuracy and the compression ratio slightly increase by up to 6%.
Lemma 1 describes how much the compression ratio increases after the proposed frequency-based post-processing.
Lemma 1.
Let $r$ and $r^{\prime}$ denote the compression ratio before and after the frequency-based post-processing. For a threshold $K$, where the tokens that appear equal or less than $K$ times are replaced with their smaller, more-frequent constituents, we have
$\frac{r^{\prime}}{r}\leq\frac{N-\sum_{k=1}^{K}\left(\lambda_{k}-\lambda_{k}w\right)}{N}.$
Here, $N$ is the number of tokens in the original MultiTok encoded data, $\lambda_{k}$ is the number of tokens that appear $k$ times in the training data, and $w$ is the training window size.
Proof.
The number of tokens before the post-processing is $N$. During the post-processing, for every $k\in\{1,\dots,K\}$, the corresponding $\lambda_{k}$ multi-word MultiTok tokens that only appear $k$ times in the training data are removed. Each of these tokens are replaced with their constituents that are either visited more than $K$ times or are single-word tokens. The number of these constituents are upper bounded by $w$ by construction. Thus, the number of tokens after the post processing is at most $N-\sum_{k=1}^{K}\left(\lambda_{k}-\lambda_{k}w\right)$, and the upper bound for the ratio of $r$ and $r^{\prime}$ is concluded. ∎
Remark.
We hypothesize that this post-processing step increases the accuracy of the downstream model, confirmed by our experimental results. In particular, when a token rarely occurs, the relation of that token with other tokens is less explored during training. Thus, the semantics can be less captured, degrading the performance.
Finally, we present the end-to-end tokenization pipeline utilizing Algorithm 1 through an example in Fig. 1. We begin by tokenizing the phrase “Alice goes to the Wonderland and Bob goes to the Wonderland Zoo”. We first find the largest existing phrase in the inputted dictionary (left) and thus tokenize “Alice” as 2. Then, we add “Alice goes” to the MultiTok dictionary (right) as a new token with index 9. We continue this process until the second mention of the phrase “goes to”. This phrase is already in the MultiTok dictionary and is thus encoded as 10 instead of tokens 1 and 5. Now, we add “goes to the” to the MultiTok dictionary with index 16. Afterwards, the tokenized samples undergo frequency analysis (purple). Particularly, we prune tokens like token 12 which were only used twice within the entire training data set $\mathcal{X}$. Thus, its instances in the dataset are replaced with the two tokens that it compresses, i.e., 7 and 4. The resulting tokenized sequence is inputted into the LLM for training.
@@ IV Simulations Results
In this section, we demonstrate the empirical performance of our tokenizer, MultiTok, with respect to the utility and efficiency goals described by Definitions 1 and 2. Further, we show various goals such as utility, accuracy, and training time can compete, and we demonstrate some parameters that can offer a good trade-off among them.
@@ IV-A Experimental Setup
The general pipeline includes tokenizing the training samples with the MultiTok, BERT, and GPT-2 tokenization schemes. These tokens were then assigned corresponding random vector embedding and inputted to train a selected model for the textual binary classification task.
Training. The datasets utilized in this section are the IMDB [26], the sst2 [27], and the AG-News [28] datasets. The first two are collections of movie reviews, each annotated with a binary sentiment. The third is a set of news pieces from different categories. Our experiments were performed on Google Colab instances (Xeon Processors @2.3Ghz with 12GB of RAM). The selected model utilizes a random vector embedding of dimension $100$ per token. This is fed to a model including a single transformer layer [29], with two attention heads, a position-wise feed-forward neural network containing two linear layers of dimensions $100$ and $200$, and a few LayerNorm [30] normalization layers of dimension $100$. Then, the nodes are flattened to an output layer for the binary classification. All models were trained for $20$ epochs utilizing batch sizes of $1000$, the Adam’s optimizer, binary cross-entropy loss, and a learning rate of $0.01$. The complete model configuration and hyperparameters can be found in our code repository https://github.com/noelkelias/multitok.
Evaluation. All our results are the averages over 20 trials with the same parameters. Specifically, we record the model’s average training loss, training accuracy, and testing accuracy per epoch. In addition, the compression ratio as well as the average area under the curve (AUC) given by the receiver operating characteristic curve (ROC) for each test were recorded and graphed. Our results for MultiTok were compared with BERT tokenization [15], GPT-2 [16] tokenization, and single-word tokenization where each unique word is considered as a token. For all these schemes, we use random vectors for embedding of each token. We note that the BERT and GPT-2 tokenizers are well-known off-the-shelve tokenizers utilized for LLM training today, and thus we targeted them as baselines in our experiments.
Our first round of experiments are focused on studying the performance MultiTok algorithm for various sizes of the training window and the testing window. In the second experiment, we apply MultiTok encoding on the BERT and GPT-2 tokens to have a cascade of two tokenization schemes, where MultiTok takes BERT or GPT-2 tokens as units of data for compression. Finally, we apply the frequency-based post processing on the MultiTok tokens for further improvements. We note that the title of rows in the table can be read as the type of tokenization, the percentage of training data to which the tokenizer is applied, and the training and testing window sizes.
@@
Figure 3: MultiTok demonstrates an average training time ($C(\epsilon)|\epsilon=0.01$) that is almost 2-3 times faster than the BERT & GPT-2 standards.
@@
BERT 1.17 0.738 0.81
GPT-2 1.26 0.733 0.80
No Compression 1 0.725 0.77
MultiTok, 100%, Max-Max 0.57 0.648 0.69
MultiTok, 100%, Max-2 0.57 0.643 0.69
@@
MultiTok, 100%, 2-2 0.6 0.701 0.75
MultiTok, 50%, Max-Max 0.8 0.661 0.72
MultiTok, 50%, Max-1 0.8 0.718 0.77
MultiTok, 50%, 2-2 0.83 0.706 0.76
MultiTok, 50%, 2-1 0.83 0.743 0.82
MultiTok, 25%, 2-1 0.93 0.728 0.78
BERT-MultiTok, 50%, 2-1 0.8 0.735 0.81
BERT-MultiTok, 25%, 2-1 0.91 0.753 0.83
GPT-2-MultiTok, 50%, 2-1 0.81 0.731 0.79
GPT-2-MultiTok, 33%, 2-1 0.88 0.745 0.82
MultiTok, 100%, 2-2, $\geq$2 0.66 0.729 0.79
MultiTok, 100%, Max-Max, $\geq$2 0.65 0.709 0.77
@@ IV-B MultiTok Tokenization
Table I shows the results for tokenization schemes with different parameters. According to this table, the best performance is obtained by the presented MultiTok tokenization when the size of training and testing windows are $2$ and $1$, respectively, and the tokenization dictionary is applied to $50\%$ of samples in the training data. We remind that the testing window size $1$ means that during the inference from the trained downstream model, we used single-word tokens, and multi-word tokens were only used during training. In this scenario, MultiTok tokenization exhibits a slightly better performance compared to the standard BERT’s performance ($0.743$ compared to $0.739$) on the IMDB dataset, while compressing the training data by $17\%$.
@@
IV-C Combining MultiTok with Existing Tokenizers
In this experiment, we used MultiTok tokenization to further compress the repetitive sequences within BERT and GPT-2 tokens. This consequently reduces the training data size while achieving similar performance to that of normal uncompressed BERT & GPT-2 tokens. We observed that we could compress the training data that is tokenized via BERT by $26\%$ and via GPT-2 by $38\%$ using MultiTok, while achieving the highest accuracies of any downstream model with $75.3\%$ and $74.5\%$ respectively. However, when we further reduced training data size, a higher compression is obtained at the cost of slight degradation in the performance. This suggests a notion of using a ‘golden ratio’ of MultiTok compression to maximize efficiency and performance. This ‘golden ratio’ may vary based on the repetition and succinctness found within different datasets.
Next, we provide an empirical training convergence analysis, which shows that MultiTok not only performs as good as the BERT & GPT-2 standards in terms of accuracy and AUC, but also directly results in a faster training process for the downstream models.
We observe in Fig. 2 that applying MultiTok tokenization to existing BERT and GPT-2 tokens not only reduces the size of the training data, but also allows the model to train faster. Specifically, the combination of MultiTok and BERT or GPT-2 tokenization achieves an average training time (as defined in Definition  2) of epoch 11 rather than epoch 20 (almost $2x$ faster) as shown in Fig. 3. Thus, these experiments suggest that MultiTok can be cascaded with existing tokenizers, like BERT and GPT-2, for smaller training data sizes, more efficient training, and improved downstream model performance.
@@ IV-D MultiTok Tokenization with Post-Processing
Our results in this experiment indicates that the post processing of MultiTok tokenization with a frequency analysis component dramatically increases the performance of the resulting tokens. We remind that without this post processing, there is a significant decrease in the model’s performance when the full capacity of MultiTok is utilized (when MultiTok is applied to the entire dataset with maximum window sizes), see Table I. However, by pruning less-utilized tokens, we were able to circumvent any impact to accuracy, as shown by the last set of experimental results. Specifically for experiment “MultiTok, 100%, 2-2, $\geq$2” (MultiTok is applied to the entire dataset with a training window size of 2), we see that by simply removing all multi-word tokens that only appear once in the training data, we get an accuracy of around $73\%$. While the training data size of the initial MultiTok experiment (MultiTok, 100%, 2-2,) increases by $6\%$, the resulting more succinct tokens provide an accuracy comparable to the BERT & GPT-2 baseline standards while still reducing the original training data size by $34\%$.
According to Fig. 2, using these post-processed MultiTok tokens can result in a comparable accuracy up to $2.5x$ faster than when using BERT or GPT-2 tokens. Most impressively, at epoch 8 in Fig. 2, the loss of this stand-alone post-processed MultiTok experiment has almost achieved its stable minimum value, being at least 0.4 lower than the training loss for the baselines at the same epoch. In addition, as shown in Fig. 3, the average training time for MultiTok based experiments is 8-12 epochs, which is significantly lower than the other counterpart tokenizers. Significantly, MultiTok has the lowest compression ratio out of the selected experiments, where the MultiTok tokens not only compressed the inputted training data by a factor of around $3$, but also helped the model learn the training data almost $2.5x$ faster. This confirms that with MultiTok tokenization, the model is able to learn the training data and existing tokens faster and more efficiently as opposed to the baseline tokenization schemes.
V Conclusion & Future Work
We proposed MultiTok, a novel multi-word tokenizer to notably compress repetitive words or phrases within a training dataset without harming the performance of downstream models. The proposed scheme can be used stand-alone and applied directly on words of text-based data, or it can be followed by an off-the-shelf tokenizer to improve efficiency and performance. Specifically, we demonstrate being able to compress training data by up to $33\%$ and train the downstream model $2.5x$ faster while maintaining comparable performance to the BERT and GPT-2 standards. A promising next step is to extend MultiTok to be used for complex NLP tasks like text generation and even other data types including videos. We hope that this paper marks the start of utilizing such data compression techniques to create more efficient LLMs of the future.